### v1

# Corresponding author(s)

Abstract
Recent advancements in slow-thinking reasoning models have shown exceptional performance in complex reasoning tasks. However, these models often exhibit overthinking-generating redundant reasoning steps for simple problems, leading to excessive computational resource usage. While current mitigation strategies uniformly reduce reasoning tokens, they risk degrading performance on challenging tasks that require extended reasoning. This paper introduces Difficulty-Adaptive Slow-Thinking (DAST), a novel framework that enables models to autonomously adjust the length of Chain-of-Thought(CoT) based on problem difficulty. We first propose a Token Length Budget (TLB) metric to quantify difficulty, then leverage length-aware reward shaping and length preference optimization to implement DAST. DAST penalizes overlong responses for simple tasks while incentivizing sufficient reasoning for complex problems. Experiments on diverse datasets and model scales demonstrate that DAST effectively mitigates overthinking (reducing token usage by over 30% on average) while preserving reasoning accuracy on complex problems.
1 Introduction
Recently, significant advancements have been made in slow-thinking reasoning models, exemplified by OpenAI’s o1[11] and DeepSeek-R1[2]. These slow-thinking reasoning models, which simulate human deep-thinking mechanisms through self-reflection, error correction, and exploration of multiple solution strategies, have demonstrated remarkable potential in complex reasoning tasks such as mathematical problem-solving[8] and programming challenges[4].
Figure 1: An Example of overthinking phenomenon from DeepSeek-R1.
However, empirical studies [1, 15] have shown that these reasoning models suffer from overthinking phenomenon. In other words, these models tend to generate redundant solutions and unnecessarily complex reasoning steps when addressing simple problems, leading to inefficient computational resource utilization. For instance, as demonstrated in Figure 1, traditional LLM can solve basic mathematic problems such as “3x + 7=22, x=?” with just 58 tokens, while reasoning model such as DeepSeek-R1[2] may consume over 1000 tokens for the same problem. This overthinking phenomenon not only significantly reduces reasoning efficiency but also causes information overload for the users.
Current approaches [1, 14] to mitigate the overthinking problem typically employ a one-size-fits-all strategy, uniformly reducing reasoning steps or token counts across all problems. Although these approaches significantly reduce the output length of slow-thinking models, they risk causing performance degradation, particularly when addressing challenging problems. Prior studies [15, 10] have demonstrated that adequate reasoning length is critical for slow-thinking models to effectively solve complex tasks. It is therefore essential to devise approaches that mitigate overthinking phenomena while maximally preserving reasoning capabilities.
This raises a fundamental question: Can we develop solutions that enable slow-thinking models to autonomously determine reasoning depth according to problem difficulty, thereby generating concise responses for simple questions while maintaining sufficiently extended chain-of-thought reasoning for complex problems? We propose a Difficulty-Adaptive Slow-Thinking (DAST) framework to address this challenge.
Our key idea is straightforward: given that tasks of varying difficulty levels inherently demand different reasoning depths, we propose to establish a mapping relationship between problem complexity and target response length. By comparing the length of the current response with the target response length, we can determine whether to apply additional rewards or penalties to the current answer. Building upon this, we construct a training objective to achieve adaptive reasoning. Specifically, we first introduce a difficulty quantification metric termed ”Token Length Budget”(TLB), which integrates both the accuracy of sampled responses and their length distributions. This metric effectively combines problem difficulty characteristics with token length information. For multiple generated responses sampled, our method applies length-aware reward shaping: Responses exceeding the TLB of simple questions receive penalty signals, while those approaching the TLB for complex problems receive positive incentives. This mechanism allows us to construct pair-wise length preference training data that inherently encodes the relationship between problem difficulty and target response length. Through follow-up preference optimization, we enable the slow-thinking model to acquire adaptive reasoning capabilities, strategically allocating more computational resources to challenging problems while maintaining efficient processing of simpler tasks. The proposed DAST method essentially establishes a learnable mapping between problem difficulty levels and corresponding target response length, achieving intelligent computation allocation during the inference stage without compromising reasoning quality.
Our main contributions are as follows:
  1. 1.
    We propose a difficulty-adaptive slow thinking (DAST) scheme, which effectively alleviates the phenomenon of overthinking while maintaining the reasoning performance, especially on difficult tasks.
  2. 2.
    We propose a novel problem difficulty quantification metric (TLB) that is applicable to many downstream tasks.
  3. 3.
    We conduct extensive validation experiments across multiple datasets with models of varying parameter scales. The results demonstrate that the proposed DAST approach effectively mitigates the overthinking problem while preserving the model’s reasoning capabilities.
2 Related Works
2.1 Problem Difficulty Evaluation
Some previous studies have primarily relied on proprietary models such as ChatGPT to assess the difficulty or complexity of data samples [7, 6]. Such methods are limited by the capabilities of LLMs used for evaluation. In reasoning scenarios such as mathematics, a more common solution is to use sampling accuracy to measure the difficulty of the problem [13, 12]. However, this approach has two limitations. First, it requires sampling a larger number of answers to ensure effectiveness. Second, for some more difficult questions, there are fewer valid answers, resulting in inadequate discriminative power. The token length budget metric we propose in this paper can effectively circumvent these two limitations.
2.2 OverThinking Solution
Currently, two studies have addressed the phenomenon of overthinking in slow-thinking models and proposed corresponding solutions. [1] introduced a method based on length contrastive optimization approach, which substantially reduces the reasoning length. [14] developed a redundant token compression method to address this issue. Nevertheless, both approaches neglect problem difficulty differentiation, leading to substantial performance degradation on complex problems. In this paper, we propose a difficulty adaptive slow thinking approach to achieve a trade-off between reasoning length and performance.
3 Methodology
In this section, we introduce our proposed DAST method in detail. Our key insight lies in enhancing existing reasoning models through length-preference training, enabling adaptive response generation with length that corresponds to problem complexity. The main challenge lies in establishing a principled relationship between response length and problem difficulty. To this end, we propose a novel reasoning Token Length Budget (TLB) metric that dynamically scales with problem complexity: simpler questions receive smaller length allocations while complex ones are allocated extended budgets. This metric not only serves as a length reference for response generation but also could be used to quantify problem difficulty.
The technical implementation of DAST involves three crucial steps: First, calibrating the initial rule-based reward scores of each response with thinking process via comparing its actual token length with the TLB of the corresponding problem. Second, constructing a pairwise length-preference training dataset based on the calibrated reward scores. Finally, employing SimPO [9] to fine-tune the original reasoning model, endowing it with adaptive reasoning capabilities. The overall framework of our proposed method is depicted in Figure 2.
Figure 2: Overview of our proposed DAST method.
3.1 Token Length Budget Definition
Our proposed Token Length Budget (TLB) metric is formally defined as:
$L_{budget}=p\cdot L_{\overline{r}}+(1-p)\cdot L_{max},$ (1)
where
$p=\frac{c}{N}$
denotes the sampling accuracy. Here, $c$ is the number of correct responses sampled from the current question, $N$ is the total number of sampled responses. $L_{\overline{r}}$ represents the average token length of correct responses, and $L_{\mathrm{max}}$ is the maximum generation length.
The higher the sampling accuracy, the closer $L_{budget}$ is to the average length of correct responses ($L_{\overline{r}}$ ), and the lower the sampling accuracy, the closer the $L_{budget}$ is to the maximum generated length. When the sampling accuracy is 0, it means that the current problem is extremely difficult, under these circumstances, the model should be encouraged to think as much as possible and generate longer COT. At this time, TLB is equal to the maximum generation length of the model. As shown in Figure 3, the average TLB exhibits strong positive correlation with problem difficulty level on the MATH training dataset, demonstrating its potential as an effective measure for quantifying problem complexity.
Figure 3: Average token budget distribution across difficulty levels ($L_{\mathrm{max}}=4096$). Results are computed using responses sampled from QwQ-32B-preview and DeepSeek-R1-Distill-Qwen-32B-preview on the MATH training set. The higher TLB for DeepSeek-R1-Distill-Qwen-32B-preview stems from its structured output format containing both reasoning chains and final answers.
3.2 Reward Score Calibration
In reasoning scenarios such as mathematics and coding, o1-like slow thinking models typically employ rule-based rewards as feedback signals for training [2, 12]. In this work, traditional rule-based rewards are calibrated by incorporating the deviation between actual response length and the TLB metric. This calibration allows the reward score to jointly capture both difficulty-aware information and length characteristics, enabling difficulty-adaptive training.
[…]
where
$\lambda=\dfrac{L_{i}-L_{budget}}{L_{budget}}$
The calibrated reward score for response $i$ is defined as Equation 2. From Figure 4, we can derive the following insights:
Figure 4: Calibrated reward function.
For a correct answer, if its length exceeds TLB, it will result in a decay of the reward. The simpler the question, the smaller the TLB. If the current generated length significantly surpasses TLB, the reward decay will be more severe. Conversely, if the generated length is below TLB, the reward will be amplified to some extent, thereby encouraging the model to generate shorter answers within corresponding TLB.
For incorrect answers, if the actual length is below TLB, it indicates insufficient reasoning. In this case, the model is encouraged to engage in more thorough thinking process and generate longer responses. The closer the length is to TLB, the greater the calibrated reward score. Once the TLB is reached, the reward score saturates.
3.3 Length Preference Data Construction
For each question $x$, $N$ candidate responses are sampled with their TLB $L_{\mathrm{budget}}^{(x)}$ computed via Equation 1. Corresponding reward scores are then derived using Equation 2. Subsequently, all responses are ranked by the reward scores to constructed contrastive pairs $(x,y_{w},y_{l})$, where $y_{w}$ strictly dominates $y_{l}$ in reward by a minimum margin $\delta$. The maximum number of contrastive pairs per question is limited to a threshold $K$.
The resultant dataset $D_{pre}$ comprises two categories of contrastive pairs: (1) pairs where both samples generate correct answers but the positive instance demonstrates significantly shorter length; (2) pairs where both responses are incorrect yet the positive sample exhibits substantially longer answer length compared to its negative counterpart.
3.4 Length Preference Training
The constructed dataset $\mathcal{D}_{\mathrm{pre}}$ enables alignment of reasoning LLMs through Simple Preference Optimization (SimPO) [9]. We chose SimPO due to its characteristic of being more sensitive in controlling answer length. The optimization objective is formulated as:
$\mathcal{L}_{\mathrm{SimPO}}(\pi_{\theta})=-E_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)\right],$ (3)
where $\beta$ and $\gamma$ are hyperparameters.
4 Experiments
4.1 Experimental Setup
Comparative experiments are conducted on two representative slow-thinking reasoning models:DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Qwen-32B. For both of them, we sample 20 answers for each question in the PRM12K MATH dataset [5] training set with sampling temperature $\tau=0.6$. The derived TLB metrics and calibrated rewards construct the SimPO training set $\mathcal{D}_{\mathrm{pre}}$ with reward margin $\delta=0.3$. Each question retains at most 4 contrastive pairs, yielding 18,231 training instances for SimPO training.
Benchmarks
Evaluation is performed on:
  • MATH-500 [5]: 500 competition mathematics problems from the MATH[3] test set
  • AIME-2024 [8]: 30 problems from the American Invitational Mathematics Examination, which tests mathematical problem solving across multiple areas
with maximum sequence length constrained to 32,768 tokens.
Baselines
Two methods from [1] addressing overthinking are reproduced:
  • SFTShortest: Supervised fine-tuning using only the shortest correct sampled response of each problem as training data
  • SimPOShortest: SimPO with contrastive instance pairs, which take the shortest correct sampled response of each problem as positive instance and the longest as negative instance
Models MATH500 AIME2024
Accuracy Tokens Accuracy Tokens
DeepSeek-R1-Distill-Qwen-7B 92.8 3506 60.0 7448
DeepSeek-R1-Distill-Qwen-7B-SFTShortest 91.8 2954 50.0 10757
DeepSeek-R1-Distill-Qwen-7B-SimPOShortest 87.8 970 33.3 2737
DeepSeek-R1-Distill-Qwen-7B-SimPODAST 92.6 2802 53.3 6337
DeepSeek-R1-Distill-Qwen-32B 94.4 3781 73.0 10955
DeepSeek-R1-Distill-Qwen-32B-SFTShortest 92.6 2401 60.0 9009
DeepSeek-R1-Distill-Qwen-32B-SimPOShortest 89.4 1181 36.7 2690
DeepSeek-R1-Distill-Qwen-32B-SimPODAST 96.4 2013 66.7 5341
Table 1: Main results
4.2 Main Results
[…]
Method Level 1 Level 2 Level 3 Level 4 Level 5
SimPOShortest 69.3% 61.0% 66.1% 70.1% 70.8%
SimPODAST 53.7% 40.1% 42.0% 41.5% 36.6%
Table 2: Comparison of the relative reduction ratios of average token lengths between SimPOShortest and SimPODAST across different difficulty levels in MATH500
We also compared the accuracy and average token length based on DeepSeek-R1-Distill-Qwen-32B on MATH500 according to difficulty level. As shown in Figures 5(a) and 5(b), SimPODAST achieved the best accuracy on Level 4 and Level 5 with significant margin against other methods, which demonstrates that it maintains its reasoning ability under complex problems. Although SimPOShortest shows the most significant reduction in response length, its reasoning capability notably declines when addressing complex problems.
Statistical analysis of output length reduction ratios in Table 2 reveals that SimPOShortest exhibits limited differentiation in length reduction across different difficulty levels. In contrast, the DAST method shows discernible adaptive capabilities, achieving approximately 53.7% length reduction at Level 1 compared to the original model, while this reduction decreases to approximately 36.6% at the most challenging Level 5. This progressive performance variation validating its difficulty-adaptive nature.
(a) Accuracy on different levels.
(b) Token length on different levels .
Figure 5: Comparative results for different difficulty levels on MATH500
Figure 6: (a) An overthinking case from DeepSeek-R1-Distill-Qwen-32B. (b) Output from DeepSeek-R1-Distill-Qwen-32B with DAST.
Figure 7: (a) An overthinking case from DeepSeek-R1-Distill-Qwen-32B. (b) Output from DeepSeek-R1-Distill-Qwen-32B with DAST.
4.3 Case Study
Figure 6 and 7 demonstrate a comparison of results for two simple problems from DeepSeek-R1-Distill-Qwen-32B. It can be observed that that the original outputs include extensive unnecessary and redundant thinking processes, while the outputs after applying DAST are more concise and focused. Figure 8 demonstrates that the DAST method maintains its reasoning capability in addressing complex problems.
Figure 8: A Level 5 case from MATH500 answered by DeepSeek-R1-Distill-Qwen-32B with DAST.
5 Conclusion
This work addresses the critical efficiency-performance dilemma in slow-thinking models through difficulty-aware reasoning adaptation. By establishing correlation between problem complexity and optimal solution length, the proposed DAST framework enables dynamic resource allocation for reasoning. Experimental validations across MATH500 and AIME2024 confirm the effectiveness of our method.